OpenStudy (anonymous):

What is the total number of electrons that can exist in the 3rd energy level? is the "3rd energy level" the 3rd period?

OpenStudy (anonymous):

2n2(square) => 2*3*3 = 18 energy level is not a period... 3rd period elements' last energy level is third...

OpenStudy (anonymous):

where did you get the formula: 2n^2?

OpenStudy (anonymous):

hard to explain..) n^2 is the number of orbitals in that energy level...and in each orbital there are 2 electrons..

OpenStudy (anonymous):

18 are the max no. of electrons in a shell by the formula 2n*n! But a shell can contain just 8 electrons! Therefore Only 8 electrons are there in the 3rd shell! Cheers!

OpenStudy (anonymous):

the question is "what is the total number of electrons that can exist..."

OpenStudy (anonymous):

yeah ! 8 only!

OpenStudy (anonymous):

wait why is it 8? using the formula the first level should have 2. the second should have 8, but technically 6 right? if you take away the 2 from the first shell. The third (the total) would be 18, but 10 in actuality, if you take away the 8 from the 2 other shells. Right? s-2 p-6 d-10 so for third it would be 10+6+2 so 18 I was a little confused at first, but i think im understanding this more :D thanks for the help guys tell me if im wrong!

OpenStudy (anonymous):

you're right) (and I think u understand it right)

OpenStudy (anonymous):

I hope that i understand it now :D i have a test on it tomorrow so wish me luck! thanks both of you for helping :)

OpenStudy (anonymous):

:) welcm!

OpenStudy (anonymous):

oh...sorry Christopher..my english is not good ..so i didnt understand you right.. actually 2n2 formula is for shell.. in the first shell there are 2 electrons (only s electrons exist, orbital is 1s - 2 electrons) in the second shell there are 8 electrons(s and p, 2s - 2 electrons, 2p - 6 electrons = 8 electrons total) in the third shell-----18 electrons( s, p and d, 3s-2 electrons, 3p-6 electrons, 3d- 10 electrons, total 18 electrons..) if we add up all of them , 2(1st)+8(2nd)+18(3rd) = 28 there may exist 28 electrons in the first 3 energy levels.(not 18->18 is the only electrons in the 3rd energy level)

OpenStudy (anonymous):

So i was right! :)